OpenStudy (anonymous):

Which point lies on the line described by the equation below? y + 8 = 4(x - 5) A. (5, -8) B. (4, -8) C. (-5, -8) D. (5, 8) E. (4, 5) F. (5, 0)

OpenStudy (bakonloverk):

you have to solve the equation

OpenStudy (anonymous):

so it is y +8= 4x - 5

OpenStudy (anonymous):

would i subtract 5 from 8?

OpenStudy (bakonloverk):

i believe so

OpenStudy (anonymous):

i got 3

OpenStudy (anonymous):

?

OpenStudy (bakonloverk):

yea its three but hold on i think thts wrong

OpenStudy (anonymous):

yeah me too

OpenStudy (bakonloverk):

sorry i cant help, im not good with two sided equations

OpenStudy (bakonloverk):

@Notamathgenius

OpenStudy (anonymous):

its ohkay

OpenStudy (notamathgenius):

Ok all you need to do is substitute the numbers for X and Y (x,y)

OpenStudy (anonymous):

so how would i do that?

OpenStudy (notamathgenius):

Ok first you take the equation y + 8 = 4(x - 5) and you take an answer choice ill use E (4, 5) You take 4 in the place of the X and 5 in the place of y and they are on both side of the = sign so if 4(4-5) = the same as 5+8 that is the answer

OpenStudy (anonymous):

yeah im offically confused

OpenStudy (anonymous):

A is

OpenStudy (bakonloverk):

just do wht he said, you do the equation y+8=4(x-5) and then put in each number from your choices and then solve them and see which ones equal the same on each side so do each of these equations A) -8+8=4(5-5) B) -8+8=4(4-5) C) -8+8=4(-5-5) D) 8+8=4(5-5) E) 5+8 =4(4-5) F) 0+8=4(5-5)

OpenStudy (anonymous):

You have to even it out and make the equation to like if you were solving a simple equation

OpenStudy (bakonloverk):

well obviously

OpenStudy (anonymous):

A is the only answer

OpenStudy (anonymous):

I have been in geometry algebra and just started algebra 2 a is the only answer that is right.

OpenStudy (anonymous):

a was the right answer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I am a hundred percent positive

OpenStudy (anonymous):

its the simple slope point formula

OpenStudy (anonymous):

can I have a metal